# Lipschitz Bandits with Stochastic Delayed Feedback

### v1

@@ 1 Introduction
Multi-armed Bandit (MAB) (auer2002finite, ) is a foundational framework for sequential decision-making under uncertainty, with widespread applications in clinical studies (villar2015multi, ), online recommendation (li2010contextual, ), and hyperparameter tuning (ding2022syndicated, ). In the standard multi-armed bandit formulation, the learner repeatedly selects an arm and immediately observes a stochastic reward drawn from an unknown distribution, using this feedback to update their policy in real time to maximize the overall cumulative reward. However, in many real-world applications, the reward signal is not always available immediately. For example, in online advertising, user engagement data such as clicks or conversions may come back to the engine minutes or even hours after an item is shown; in clinical trials, the effect of a treatment on a patient’s health may be delayed and take weeks to observe. This motivates the study of bandit problems with delayed feedback, where the learner must make decisions without knowing when past rewards will arrive, introducing additional uncertainty not only in the stochastic reward values but also in the timing of their observations. To deal with this real-world challenge, recent progress has extended classical bandit algorithms to accommodate stochastic delays under multiple bandit frameworks, including MABs (joulani2013online, ; vernade2017stochastic, ), linear bandits (vernade2020linear, ) and kernelized bandits (vakili2023delayed, ). These approaches often rely on modifying confidence intervals or restructuring update schedules to compensate for missing feedback, and they yield robust performance under various stochastic delay models.
@@
While delayed feedback has been extensively studied in the context of stochastic bandit problems, most existing literature focuses on simpler settings with discrete action spaces, such as the traditional MABs and contextual linear bandits. In contrast, there is little understanding of how to mitigate the effects of delayed observations in the more challenging Lipschitz bandit framework (kleinberg2008multi, ), where the action space is a continuous metric space and the expected reward function satisfies a Lipschitz condition. Lipschitz bandits provide a more flexible and expressive framework for a wide range of real-world applications, including hyperparameter optimization (kang2023online, ), dynamic pricing, and auction design (slivkins2019introduction, ). State-of-the-art algorithms for Lipschitz bandits, such as the Zooming algorithm (kleinberg2019bandits, ), leverage adaptive partitioning and activation schemes to concentrate exploration on more promising regions, achieving optimal regret bounds in the delay-free setting. However, these methods fundamentally rely on prompt reward feedback to refine partitions and update confidence estimates. In the presence of stochastic delays, they may struggle to accurately track arm performance and quality, leading to degraded performance both theoretically and empirically. Moreover, the continuous nature of the arm space amplifies the impact of missing or stale feedback, as each sampled point represents a neighborhood region whose estimation depends on delayed observations. To the best of our knowledge, the Lipschitz bandit problem with stochastic delayed feedback has not been previously explored, and poses distinct challenges that require novel algorithmic designs due to the dual complexity of delayed observations and continuous action spaces.
@@
  • For bounded delays, we extend the classic zooming algorithm in kleinberg2008multi to a delay-aware setting and attains $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}+\tau_{\max}T^{\frac{d_{z}}{d_{z}+2}}\right)$ regret bound, recovering the optimal regret bound of Lipschitz bandit (kleinberg2019bandits, ), with an additional term scales with the maximal delay $\tau_{\max}$.
@@ 2 Related work
The Lipschitz bandit problem, also known as continuum-armed bandits, was first introduced in agrawal1995continuum . The problem was then extended to the general metric space in kleinberg2008multi . A common approach to solving Lipschitz bandit problems is to discretize the continuous arm spac, either uniformly or adaptively, and then apply standard finite-armed bandit algorithms to deal with the exploration-exploitation tradeoff. This reduction enables the use of strategies such as the Upper Confidence Bound (UCB) method (kleinberg2019bandits, ), Thompson Sampling (TS) (vernade2020linear, ) and elimination-based techniques (feng2022lipschitz, ). Beyond the standard stochastic setting, kang2023online ; nguyen2025non studied the non-stationary Lipschitz bandit problem, and podimata2021adaptive adapted the discretization-based framework to handle adversarial rewards case. Moreover, kang2023robust investigated the Lipschitz bandits with adversarial corruptions to bridge the gap between purely stochastic and fully adversarial regimes. In addition, several extensions of the Lipschitz bandit framework have also been explored. For instance, bandits with heavy-tailed rewards are studied in lu2019optimal , while contextual Lipschitz bandits have been addressed in slivkins2019introduction ; krishnamurthy2020contextual , and quantum variants of Lipschitz bandits have been recently investigated in yi2025quantum .
@@
The stochastic multi-armed bandit (MAB) problem with randomized delays has been extensively studied in the literature. joulani2013online analyzed how delays affect regret in online learning through BOLD and QPM-D settings and present a direct modification of the UCB1 algorithm, showing that delays lead to multiplicative regret increases in adversarial settings and additive increases in stochastic ones. vernade2017stochastic investigated delayed conversions in stochastic MABs, allowing for censored feedback under the assumption of known delay distributions. Variants involving delayed, aggregated, and anonymous feedback were addressed in pike2018bandits , where a phase-based elimination algorithm based on the Improved UCB algorithm by Auer2010UCBRI was proposed. Unrestricted and potentially unbounded delay distributions have been explored in both reward-independent and reward-dependent settings in lancewicki2021stochastic by adapting quantile function of the delay distribution into the regret bound. Recent work have extended delayed feedback setting to a variety of bandit problems, such as linear bandits (vernade2020linear, ), contextual MAB (arya2020randomized, ), generalized (contextual) linear bandits (zhou2019learning, ; howson2023delayed, ), kernel bandits (vakili2023delayed, ), Bayesian optimization (verma2022bayesian, ), dueling bandits (yi2024biased, ), adversarial bandits (zimmert2020optimal, ) and best-of-both-worlds algorithms (masoudian2024best, ). However, the study of Lipschitz bandits under delayed feedback remains an unaddressed open problem due to the unique challenges posed by the continuous arm space, where existing methods for handling delay fail.
@@ 3 Problem Settings and Preliminaries
We study the problem of Lipschitz bandit with delayed feedback. Formally, a Lipschitz bandit problem is defined on a triplet $(\mathcal{A},\mathcal{D},\mu)$, where the action space $\mathcal{A}$ is a compact doubling metric space equipped with metric $\mathcal{D}$, both known to the agent. The unknown expected reward function $\mu:\mathcal{A}\to[0,1]$ is a $1$-Lipschitz function defined on $\mathcal{A}$ w.r.t. the metric $\mathcal{D}$, i.e. $|\mu(x_{1})-\mu(x_{2})|\leq\mathcal{D}(x_{1},x_{2}),\quad\forall x_{1},x_{2}\in\mathcal{A}.$ We refer to the triple ($\mathcal{A},\mathcal{D},\mu$) as an instance of the Lipschitz bandit problem. And without loss of generality, we assume that the diameter of $(\mathcal{A},\mathcal{D})$ is bounded by $1$, which is a standard assumption as in kleinberg2019bandits .
@@
An important pair of concepts in Lipschitz bandits for a problem instance ($\mathcal{A},\mathcal{D},\mu$) are the covering dimension $d$ and zooming dimension $d_{z}$ (kleinberg2008multi, ). Let $B(x,r):=\{x^{\prime}\in\mathcal{A}:\mathcal{D}(x,x^{\prime})\leq r\}$ denote a closed ball centered at $x$ with radius $r$ in ($\mathcal{A},\mathcal{D}$). Let $\mathcal{S}$ be a subset of $\mathcal{A}$, then a subset $\mathcal{C}$ of $\mathcal{A}$ is a $r$-covering of $\mathcal{S}$ if $\mathcal{S}\subseteq\bigcup_{x\in\mathcal{C}}B(x,r)$. The $r$-covering number $N_{c}(r)$ of metric space ($\mathcal{A},\mathcal{D}$) is defined as the least number of balls with radius no more than $r$ required to completely cover $(\mathcal{A},\mathcal{D})$, with possible overlaps between the balls. The $c$-covering dimension $d$ is defined as the smallest $d$ such that for every $r>0$ we require only $O(r^{-d})$ balls with radius no more than $r$ to cover the metric space $(\mathcal{A},\mathcal{D})$:
$d=\inf\left\{n\geq 0:\exists c>0,\forall r>0,N_{c}(r)\leq cr^{-n}\right\}.$
On the other hand, the $r$-zooming number $N_{z}(r)$ and the zooming dimension $d_{z}$ depend not only on the metric space $(\mathcal{A},\mathcal{D})$ but also on the underlying payoff function $\mu$. Define the $r$-optimal region as $\{x\in\mathcal{A}:\Delta(x)\leq r\}$, where $\Delta(x)=\mu^{*}-\mu(x)$ denotes the suboptimality gap at point $x$. The $r$-zooming number $N_{z}(r)$ is the minimal number of balls of radius at most $r/16$ required to cover the $r$-optimal region. The $c$-zooming dimension $d_{z}$ is the smallest value $d$ such that, for all $r\in(0,1]$, the $r$-optimal region can be covered by $O(r^{-d})$ balls of radius at most $r/16$:
$d_{z}=\inf\left\{n\geq 0:\exists c>0,\forall r\in(0,1],N_{z}(r)\leq cr^{-n}\right\}.$
@@ 4 Delay with Bounded Support: Delayed Zooming Algorithm
We first consider the case of bounded delays. Specifically, we assume there exists a positive integer $\tau_{\max}$ such that the delay $\tau_{t}$ is supported on $\{0,1,\dots,\tau_{\max}\}$. This implies that feedback is always eventually observed and never missing, with a maximum delay of $\tau_{\max}$ rounds. To address this setting, we propose a delay-aware variant of the zooming algorithm, called the Delayed Zooming algorithm, which maintains the optimal regret rate of the delay-free setting up to an additive term that scales with the maximal delay. The algorithm achieves a regret bound of order $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}+\tau_{\max}T^{\frac{d_{z}}{d_{z}+2}}\right)$.
@@
Similar to the classic zooming algorithm, our delayed variant leverages the UCB principle and adaptive discretization (slivkins2019introduction, ). It maintains an active set of arms such that all arms in $\mathcal{A}$ are covered by the confidence balls of active arms. If an arm is not covered, it is activated. This mechanism allows the algorithm to adaptively focus on high-reward regions. In each round, the agent selects the active arm with the highest index $I_{t}(x)=\mu_{t}(x)+2r_{t}(x)$ (breaking ties arbitrarily) and pulls it. The confidence radius $r_{t}(x)$ is defined below based on the confidence level $\delta$.
@@
In the classic zooming algorithm, the agent updates the empirical mean reward and the corresponding confidence radius immediately after pulling an arm. However, in the presence of delay, these updates cannot be performed in real time, as the agent lacks access to both unobserved rewards and the number of pending pulls for each arm. Inspired by the idea of Delayed-UCB1 (joulani2013online, ), we modify the classic zooming algorithm by replacing $n_{t}(x)$, defined as the number of times arm $x$ has been pulled, with $v_{t}(x)$, the number of observed rewards from arm $x$, and compute the empirical mean using only the available feedback. While this modification appears direct and natural given the prior literature on bandits with delayed feedback (vernade2017stochastic, ), we highlight that the analysis is far from trivial and significantly different from that of multi-armed bandit with delays and the classic zooming algorithm. Specifically, the regret analysis of the classic zooming algorithm relies on a key result that bounds the sub-optimality gap by the confidence radius, i.e. $\Delta(x)\leq 3r_{t}(x)$, which ensures that suboptimal arms are not pulled too frequently. This bound holds in the delay-free setting because $r_{t}(x)$ only changes when arm $x$ is played. Under delayed feedback, however, $r_{t}(x)$ can shrink over time even when the arm is not being pulled, due to the arrival of delayed observations. This can cause the confidence radius to decrease too rapidly. In contrast, the extension from UCB1 to Delayed-UCB1 does not have this complication since its analysis on MAB is simpler and does not depend on such a suboptimality gap inequality. To address this challenge, we develop a “lazy update” mechanism showing that, under bounded delays, a similar guarantee still holds: the sub-optimality gap is bounded by $6r_{t}(x)$. Establishing this result is not a straightforward extension of the original proof and requires careful treatment of the delayed information flow, yet it notably preserves the theoretical guarantee of the classic zooming algorithm up to a constant factor.
@@
The complete description is given in Algorithm 1. As in the classic zooming algorithm, our delayed zooming algorithm also requires a covering oracle (i.e. line 5 of Algorithm 1) which takes a finite collection of closed balls and either declares that they cover all arms in the metric space or outputs an uncovered arm to apply the activation rule (kleinberg2019bandits, ). Note that if the delays are zero, this algorithm reduces to the original zooming algorithm.
@@
In the following Theorem 1, we show that the Algorithm 1 enjoys the same regret rate compared to its non-delayed version, with an additive penalty depending on the expected delay.
@@
When there is no delay, i.e. $\tau_{max}=0$, the regret bound reduces to $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}\right)$, which is exactly the regret bound of Lipschitz bandits. When the action space is finite, i.e. $d_{z}=0$, the regret bound reduces to $O(\sqrt{cT\log T}+c\tau_{\max})$, with $c$ being the number of arms, which is the regret bound of $c$-armed bandits with bounded delays in joulani2013online .

### v2

@@ 1 Introduction
Multi-armed Bandit (MAB) (Auer et al., 2002) is a foundational framework for sequential decision-making under uncertainty, with widespread applications in clinical studies (Villar et al., 2015), online recommendation (Li et al., 2010), and hyperparameter tuning (Ding et al., 2022). In the standard multi-armed bandit formulation, the learner repeatedly selects an arm and immediately observes a stochastic reward drawn from an unknown distribution, using this feedback to update their policy in real time to maximize the overall cumulative reward. However, in many real-world applications, the reward signal is not always available immediately. For example, in online advertising, user engagement data such as clicks or conversions may come back to the engine minutes or even hours after an item is shown; in clinical trials, the effect of a treatment on a patient’s health may be delayed and take weeks to observe. This motivates the study of bandit problems with delayed feedback, where the learner must make decisions without knowing when past rewards will arrive, introducing additional uncertainty not only in the stochastic reward values but also in the timing of their observations. To deal with this real-world challenge, recent progress has extended classical bandit algorithms to accommodate stochastic delays under multiple bandit frameworks, including MABs (Joulani et al., 2013; Vernade et al., 2017), linear bandits (Vernade et al., 2020) and kernelized bandits (Vakili et al., 2023). These approaches often rely on modifying confidence intervals or restructuring update schedules to compensate for missing feedback, and they yield robust performance under various stochastic delay models.
@@
While delayed feedback has been extensively studied in the context of stochastic bandit problems, most existing literature focuses on simpler settings with discrete action spaces, such as the traditional MABs and contextual linear bandits. In contrast, there is little understanding of how to mitigate the effects of delayed observations in the more challenging Lipschitz bandit framework (Kleinberg et al., 2008), where the action space is a continuous metric space and the expected reward function satisfies a Lipschitz condition. Lipschitz bandits provide a more flexible and expressive framework for a wide range of real-world applications, including hyperparameter optimization (Kang et al., 2024), dynamic pricing, and auction design (Slivkins and others, 2019). State-of-the-art algorithms for Lipschitz bandits, such as the Zooming algorithm (Kleinberg et al., 2019), leverage adaptive partitioning and activation schemes to concentrate exploration on more promising regions, achieving optimal regret bounds in the delay-free setting. However, these methods fundamentally rely on prompt reward feedback to refine partitions and update confidence estimates. In the presence of stochastic delays, they may struggle to accurately track arm performance and quality, leading to degraded performance both theoretically and empirically. Moreover, the continuous nature of the arm space amplifies the impact of missing or stale feedback, as each sampled point represents a neighborhood region whose estimation depends on delayed observations. To the best of our knowledge, the Lipschitz bandit problem with stochastic delayed feedback has not been previously explored, and poses distinct challenges that require novel algorithmic designs due to the dual complexity of delayed observations and continuous action spaces.
@@
  • For bounded delays, we extend the classic zooming algorithm in Kleinberg et al. (2008) to a delay-aware setting and attains $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}+\tau_{\max}T^{\frac{d_{z}}{d_{z}+2}}\right)$ regret bound ($d_{z}$ is the zooming dimension), recovering the optimal regret bound of Lipschitz bandit (Kleinberg et al., 2019), with an additional term scales with the maximal delay $\tau_{\max}$.
@@ 2 Related work
The Lipschitz bandit problem, also known as continuum-armed bandits, was first introduced in Agrawal (1995). The problem was then extended to the general metric space in Kleinberg et al. (2008). A common approach to solving Lipschitz bandit problems is to discretize the continuous arm spac, either uniformly or adaptively, and then apply standard finite-armed bandit algorithms to deal with the exploration-exploitation tradeoff. This reduction enables the use of strategies such as the Upper Confidence Bound (UCB) method (Kleinberg et al., 2019), Thompson Sampling (TS) (Vernade et al., 2020) and elimination-based techniques (Feng et al., 2022). Beyond the standard stochastic setting, Kang et al. (2024); Nguyen et al. (2025) studied the non-stationary Lipschitz bandit problem, and Podimata and Slivkins (2021) adapted the discretization-based framework to handle adversarial rewards case. Moreover, Kang et al. (2023) investigated the Lipschitz bandits with adversarial corruptions to bridge the gap between purely stochastic and fully adversarial regimes. In addition, several extensions of the Lipschitz bandit framework have also been explored. For instance, bandits with heavy-tailed rewards are studied in Lu et al. (2019), while contextual Lipschitz bandits have been addressed in Slivkins and others (2019); Krishnamurthy et al. (2020), and quantum variants of Lipschitz bandits have been recently investigated in Yi et al. (2025).
@@
The stochastic multi-armed bandit (MAB) problem with randomized delays has been extensively studied in the literature. Joulani et al. (2013) analyzed how delays affect regret in online learning through partial monitoring settings and present a direct modification of the UCB1 algorithm, showing that delays lead to multiplicative regret increases in adversarial settings and additive increases in stochastic ones. Vernade et al. (2017) investigated delayed conversions in stochastic MABs, allowing for censored feedback under the assumption of known delay distributions. Variants involving delayed, aggregated, and anonymous feedback were addressed in Pike-Burke et al. (2018), where a phase-based elimination algorithm based on the Improved UCB algorithm by Auer and Ortner (2010) was proposed. Unrestricted and potentially unbounded delay distributions have been explored in both reward-independent and reward-dependent settings in Lancewicki et al. (2021) by adapting quantile function of the delay distribution into the regret bound. Recent work have extended delayed feedback setting to a variety of bandit problems, such as linear bandits (Vernade et al., 2020), contextual MAB (Arya and Yang, 2020), generalized (contextual) linear bandits (Zhou et al., 2019; Howson et al., 2023), adaptivity and confounding in MAB (Qin and Russo, 2022), delayed payoff in MAB (Schlisselberg et al., 2025), Thompson Sampling (Wu and Wager, 2022; McDonald et al., 2023), kernel bandits (Vakili et al., 2023), Bayesian optimization (Verma et al., 2022), dueling bandits (Yi et al., 2024), adversarial bandits (Zimmert and Seldin, 2020) and best-of-both-worlds algorithms (Masoudian et al., 2024). However, the study of Lipschitz bandits under delayed feedback remains an unaddressed open problem due to the unique challenges posed by the continuous arm space, where existing methods for handling delay fail.
@@ 3 Problem Settings and Preliminaries
We study the problem of Lipschitz bandit with delayed feedback. Formally, a Lipschitz bandit problem is defined on a triplet $(\mathcal{A},\mathcal{D},\mu)$, where the action space $\mathcal{A}$ is a compact doubling metric space equipped with metric $\mathcal{D}$, both known to the agent. The unknown expected reward function $\mu:\mathcal{A}\to[0,1]$ is a $1$-Lipschitz function defined on $\mathcal{A}$ w.r.t. the metric $\mathcal{D}$, i.e. $|\mu(x_{1})-\mu(x_{2})|\leq\mathcal{D}(x_{1},x_{2}),\quad\forall x_{1},x_{2}\in\mathcal{A}.$ We refer to the triple ($\mathcal{A},\mathcal{D},\mu$) as an instance of the Lipschitz bandit problem. And without loss of generality, we assume that the diameter of $(\mathcal{A},\mathcal{D})$ is bounded by $1$, which is a standard assumption as in Kleinberg et al. (2019).
@@
An important pair of concepts in Lipschitz bandits for a problem instance ($\mathcal{A},\mathcal{D},\mu$) are the covering dimension $d$ and zooming dimension $d_{z}$ (Kleinberg et al., 2008). Let $B(x,r):=\{x^{\prime}\in\mathcal{A}:\mathcal{D}(x,x^{\prime})\leq r\}$ denote a closed ball centered at $x$ with radius $r$ in ($\mathcal{A},\mathcal{D}$). Let $\mathcal{S}$ be a subset of $\mathcal{A}$, then a subset $\mathcal{C}$ of $\mathcal{A}$ is a $r$-covering of $\mathcal{S}$ if $\mathcal{S}\subseteq\bigcup_{x\in\mathcal{C}}B(x,r)$. The $r$-covering number $N_{\texttt{cov}}(r)$ of metric space ($\mathcal{A},\mathcal{D}$) is defined as the least number of balls with radius no more than $r$ required to completely cover $(\mathcal{A},\mathcal{D})$, with possible overlaps between the balls. The $c$-covering dimension $d$ is defined as the smallest $d$ such that for every $r>0$ we require only $O(r^{-d})$ balls with radius no more than $r$ to cover the metric space $(\mathcal{A},\mathcal{D})$:
$d=\inf\left\{n\geq 0:\exists c>0,\forall r>0,N_{\texttt{cov}}(r)\leq cr^{-n}\right\}.$
On the other hand, the $r$-zooming number $N_{\texttt{zom}}(r)$ and the zooming dimension $d_{z}$ depend not only on the metric space $(\mathcal{A},\mathcal{D})$ but also on the underlying payoff function $\mu$. Define the $r$-optimal region as $\{x\in\mathcal{A}:\Delta(x)\leq r\}$, where $\Delta(x)=\mu^{*}-\mu(x)$ denotes the suboptimality gap at point $x$. The $r$-zooming number $N_{z}(r)$ is the minimal number of balls of radius at most $r/16$ required to cover the $r$-optimal region. The $c$-zooming dimension $d_{z}$ is the smallest value $d$ such that, for all $r\in(0,1]$, the $r$-optimal region can be covered by $O(r^{-d})$ balls of radius at most $r/16$:
$d_{z}=\inf\left\{n\geq 0:\exists c>0,\forall r\in(0,1],N_{\texttt{zom}}(r)\leq cr^{-n}\right\}.$
@@ 4 Delay with Bounded Support: Delayed Zooming Algorithm
We first consider the case of bounded delays. Specifically, we assume there exists a positive integer $\tau_{\max}$ such that the delay $\tau_{t}$ is supported on $\{0,1,\dots,\tau_{\max}\}$. This implies that feedback is always eventually observed and never missing for all rounds prior to $t=T-\tau_{\max}$, with a maximum delay of $\tau_{\max}$ rounds. To address this setting, we propose a delay-aware variant of the zooming algorithm, called the Delayed Zooming algorithm, which maintains the optimal regret rate of the delay-free setting up to an additive term that scales with the maximal delay. The algorithm achieves a regret bound of order $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}+\tau_{\max}T^{\frac{d_{z}}{d_{z}+2}}\right)$.
@@
Similar to the classic zooming algorithm, our delayed variant leverages the UCB principle and adaptive discretization (Slivkins and others, 2019). It maintains an active set of arms such that all arms in $\mathcal{A}$ are covered by the confidence balls of active arms. If an arm is not covered, it is activated. This mechanism allows the algorithm to adaptively focus on high-reward regions. In each round, the agent selects the active arm with the highest index $I_{t}(x)=\mu_{t}(x)+2r_{t}(x)$ (breaking ties arbitrarily) and pulls it. The confidence radius $r_{t}(x)$ is defined below based on the confidence level $\delta$.
@@
In the classic zooming algorithm, the agent updates the empirical mean reward and the corresponding confidence radius immediately after pulling an arm. However, in the presence of delay, these updates cannot be performed in real time, as the agent lacks access to both unobserved rewards and the number of pending pulls for each arm. Inspired by the idea of Delayed-UCB1 (Joulani et al., 2013), we modify the classic zooming algorithm by replacing $n_{t}(x)$, defined as the number of times arm $x$ has been pulled, with $v_{t}(x)$, the number of observed rewards from arm $x$, and compute the empirical mean using only the available feedback. While this modification appears direct and natural given the prior literature on bandits with delayed feedback (Vernade et al., 2017), we highlight that the analysis is far from trivial and significantly different from that of multi-armed bandit with delays and the classic zooming algorithm. Specifically, the regret analysis of the classic zooming algorithm relies on a key result that bounds the sub-optimality gap by the confidence radius, i.e. $\Delta(x)\leq 3r_{t}(x)$, which ensures that suboptimal arms are not pulled too frequently. This bound holds in the delay-free setting because $r_{t}(x)$ only changes when arm $x$ is played. Under delayed feedback, however, $r_{t}(x)$ can shrink over time even when the arm is not being pulled, due to the arrival of delayed observations. This can cause the confidence radius to decrease too rapidly. In contrast, the extension from UCB1 to Delayed-UCB1 does not have this complication since its analysis on MAB is simpler and does not depend on such a suboptimality gap inequality. To address this challenge, we develop a “lazy update” mechanism showing that, under bounded delays, a similar guarantee still holds: the sub-optimality gap is bounded by $6r_{t}(x)$. Establishing this result is not a straightforward extension of the original proof and requires careful treatment of the delayed information flow, yet it notably preserves the theoretical guarantee of the classic zooming algorithm up to a constant factor.
@@
The complete description is given in Algorithm 1. As in the classic zooming algorithm, our delayed zooming algorithm also requires a covering oracle (i.e. line 5 of Algorithm 1) which takes a finite collection of closed balls and either declares that they cover all arms in the metric space or outputs an uncovered arm to apply the activation rule (Kleinberg et al., 2019). Note that if the delays are zero, this algorithm reduces to the original zooming algorithm.
@@
In the following Theorem 1, we show that the Algorithm 1 enjoys the same regret rate compared to its non-delayed version, with an additive penalty depending on the maximum delay.
@@
When there is no delay, i.e. $\tau_{max}=0$, the regret bound reduces to $\tilde{O}\left(T^{\frac{d_{z}+1}{d_{z}+2}}\right)$, which is exactly the regret bound of Lipschitz bandits. When the action space is finite, i.e. $d_{z}=0$, the regret bound reduces to $O(\sqrt{cT\log T}+c\tau_{\max})$, with $c$ being the number of arms, which is the regret bound of $c$-armed bandits with bounded delays in Joulani et al. (2013).